# Bibliography1.bib

### v2

@@ 3.2. Boundedness and indecomposability of minimizers
We remark that Theorem 3.6 proves boundedness of minimizers without requiring decreasing properties of the Riesz-type kernel, but only infinitesimality and symmetry.
@@
$\begin{split}\mathscr{R}(\tilde{E}_{t})-\mathscr{R}(E)&=\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x+\int_{\tilde{E}_{t}\cap E}\int_{\tilde{E}_{t}\setminus E}g(y-x)\,{\rm d}y\,{\rm d}x\\ &\quad-\int_{E\setminus\tilde{E}_{t}}\int_{E}g(y-x)\,{\rm d}y\,{\rm d}x-\int_{E\cap\tilde{E}_{t}}\int_{E\setminus\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x\\ &=\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x+\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}\cap E}g(x-y)\,{\rm d}y\,{\rm d}x\\ &\quad-\int_{E\setminus\tilde{E}_{t}}\int_{E}g(y-x)\,{\rm d}y\,{\rm d}x-\int_{E\setminus\tilde{E}_{t}}\int_{E\cap\tilde{E}_{t}}g(x-y)\,{\rm d}y\,{\rm d}x\\ &=\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x+\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}\cap E}g(y-x)\,{\rm d}y\,{\rm d}x\\ &\quad-\int_{E\setminus\tilde{E}_{t}}\int_{E}g(y-x)\,{\rm d}y\,{\rm d}x-\int_{E\setminus\tilde{E}_{t}}\int_{E\cap\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x\\ &\leq\int_{\tilde{E}_{t}\Delta E}\left(\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y+\int_{E}g(y-x)\,{\rm d}y\right)\,{\rm d}x\end{split}$
@@ 4. Nonexistence of minimizers for large masses
Let $g$ be $\mathscr{R}$-admissible, $q$-decreasing and infinitesimal and let $G$ be $\mathscr{G}$-admissible. There exists $\varepsilon>0$, depending on $n$, $\lambda$, $g$, $G$ and $q$, such that the following holds. Let $F\subset\mathbb{R}^{n}\setminus H$ be a set of finite perimeter and assume there exist two sets of finite perimeter $F_{1}$, $F_{2}\subset F$ such that $|F_{1}|,$ $|F_{2}|>0$, $|F_{1}\cap F_{2}|=0$, $|F\setminus(F_{1}\cup F_{2})|=0$ and
@@
Let $g$ be $\mathscr{R}$-admissible, $q$-decreasing and infinitesimal, and let $G$ be $\mathscr{G}$-admissible. Then there exists $c=c(n,\lambda,g,G,q)>0$ such that the following holds. Let $E\subset\mathbb{R}^{n}\setminus H$ be a minimizer of $\mathscr{F}^{\lambda}$ with $|E|=m$, $m>0$. Then for almost every $x\in E$ there holds
@@
We remark that the density estimate in Lemma 4.2 is more precise than the one provided in Section 3.2. Indeed, in Lemma 3.11 $K$ and $r_{0}$ depend on the minimizer, and consequently $c$ in Theorem 3.9 also inherits this dependence. At the same time, Lemma 4.2 requires $g$ to be $q$-decreasing, which is not required in Lemma 3.11.
@@ 4.1. Absence of holes in minimizers
Let $g$ be $\mathscr{R}$-admissible, 00-decreasing, infinitesimal and symmetric and let $G$ be $\mathscr{G}$-admissible and coercive. There exists $\bar{m}>0$, depending on $n$, $\lambda$, $g$ and $G$, such that, for all $m\in(0,\bar{m})$, every minimizer $E$ of $\mathscr{F}^{\lambda}$ with $|E|=m$ has the following property. There is no set $F\subset\mathbb{R}^{n}\setminus(H\cup E)$ with $|F|>0$ such that
@@
Let $g$ be $\mathscr{R}$-admissible, $q$-decreasing, infinitesimal and symmetric and let $G$ be $\mathscr{G}$-admissible and coercive. There exist $\bar{m}>0$ and $\bar{T}>0$, depending on $n$, $\lambda$, $g$, $G$, $q$ such that, for all $m\in(0,\bar{m})$, every volume constrained minimizer $E$ of $\mathscr{F}^{\lambda}$ with $|E|=m$ satisfies
@@
We remark that Lemma 4.8 is a stronger result than Lemma 3.10. Indeed, the bound in Lemma 4.8 does not depend on the minimizer. At the same time, Lemma 3.10 does not require that $g$ is $q$-decreasing and $G$ is coercive.
@@ 5. Generalized minimizers
Let $g$ be $\mathscr{R}$-admissible and $q$-decreasing and let $G$ be $\mathscr{G}$-admissible. For every $\varepsilon_{1},\varepsilon_{2}>0$ there exists a minimizer of ~$\tilde{\mathscr{F}}^{\lambda}_{\varepsilon}$ in the class
@@
Let $E\subset\mathbb{R}^{n}\setminus H$ be a set of finite perimeter. Let $g$ be $\mathscr{R}$-admissible and $q$-decreasing and let $G$ be $\mathscr{G}$-admissible. If $\alpha>1$, then
@@
(5.5) $\sigma(b^{-})\leq\frac{1}{4}n|B^{\lambda}|^{\frac{1}{n}}m^{\frac{n-1}{n}}.$
@@
Let $g$ be $\mathscr{R}$-admissible and $q$-decreasing and let $G$ be $\mathscr{G}$-admissible. Let $\varepsilon_{1},\varepsilon_{2}>0$. For every $m\in(0,\infty)$ there exist $R>0$ and $\bar{h}\in\mathbb{N}$, depending on $n$, $\lambda$, $m$, $\varepsilon_{1}$, $\varepsilon_{2}$, $g$, $G$ and $q$, such that
@@
(5.12) $m_{1}\geq\left(\frac{n|B^{\lambda}|^{\frac{n+1}{n}}}{4K}\right)^{n}.$
@@
(5.13) $\alpha=\left(\frac{|B^{\lambda}|}{|B^{\lambda}|-\sum_{j>h^{\prime}}m_{j}}\right)^{\frac{1}{n}}\leq 1+c_{1}\sum_{j=h^{\prime}+1}^{h}m_{j},$
@@
Note that ~$\tilde{\mathscr{F}}^{\lambda}(\tilde{F})=\bar{\varepsilon}^{n-1}\tilde{\mathscr{F}}_{\bar{\varepsilon}}^{\lambda}(F)$ and that, if $g$ is $q$-decreasing, also $\tilde{g}$ is $q$-decreasing (see Remark 5.2). By Theorem 5.3 there exists $E\subset\mathbb{R}^{n}\setminus H$ with $|E|=|B^{\lambda}|$ which minimizes ~$\tilde{\mathscr{F}}^{\lambda}_{\bar{\varepsilon}}$. Then the set $\tilde{E}:=\bar{\varepsilon}E$ minimizes ~$\tilde{\mathscr{F}}^{\lambda}$ among sets with volume $m$ and Theorem 1.3 easily follows. ∎

### v3

@@ 3.2. Boundedness and indecomposability of minimizers
We remark that Theorem 3.8 proves boundedness of minimizers without requiring growing properties of the Riesz-type kernel, but only infinitesimality and symmetry.
@@
$\begin{split}\mathscr{R}(\tilde{E}_{t})-\mathscr{R}(E)&=\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x+\int_{\tilde{E}_{t}\cap E}\int_{\tilde{E}_{t}\setminus E}g(y-x)\,{\rm d}y\,{\rm d}x\\ &\quad-\int_{E\setminus\tilde{E}_{t}}\int_{E}g(y-x)\,{\rm d}y\,{\rm d}x-\int_{E\cap\tilde{E}_{t}}\int_{E\setminus\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x\\ &=\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x+\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}\cap E}g(x-y)\,{\rm d}y\,{\rm d}x\\ &\quad-\int_{E\setminus\tilde{E}_{t}}\int_{E}g(y-x)\,{\rm d}y\,{\rm d}x-\int_{E\setminus\tilde{E}_{t}}\int_{E\cap\tilde{E}_{t}}g(x-y)\,{\rm d}y\,{\rm d}x\end{split}$
$\begin{split}&=\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x+\int_{\tilde{E}_{t}\setminus E}\int_{\tilde{E}_{t}\cap E}g(y-x)\,{\rm d}y\,{\rm d}x\\ &\quad-\int_{E\setminus\tilde{E}_{t}}\int_{E}g(y-x)\,{\rm d}y\,{\rm d}x-\int_{E\setminus\tilde{E}_{t}}\int_{E\cap\tilde{E}_{t}}g(y-x)\,{\rm d}y\,{\rm d}x\\ &\leq\int_{\tilde{E}_{t}\Delta E}\left(\int_{\tilde{E}_{t}}g(y-x)\,{\rm d}y+\int_{E}g(y-x)\,{\rm d}y\right)\,{\rm d}x\end{split}$
@@ 4. Nonexistence of minimizers for large masses
Let $g$ be $\mathscr{R}$-admissible, $q$-growing and infinitesimal and let $G$ be $\mathscr{G}$-admissible. There exists $\varepsilon>0$, depending on $n$, $\lambda$, $g$, $G$ and $q$, such that the following holds. Let $F\subset\mathbb{R}^{n}\setminus H$ be a set of finite perimeter and assume there exist two sets of finite perimeter $F_{1}$, $F_{2}\subset F$ such that $|F_{1}|,$ $|F_{2}|>0$, $|F_{1}\cap F_{2}|=0$, $|F\setminus(F_{1}\cup F_{2})|=0$ and
@@
Let $g$ be $\mathscr{R}$-admissible, $q$-growing and infinitesimal, and let $G$ be $\mathscr{G}$-admissible. Then there exists $c=c(n,\lambda,g,G,q)>0$ such that the following holds. Let $E\subset\mathbb{R}^{n}\setminus H$ be a minimizer of $\mathscr{F}^{\lambda}$ with $|E|=m$, $m>0$. Then for almost every $x\in E$ there holds
@@
We remark that the density estimate in Lemma 4.2 is more precise than the one provided in Section 3.2. Indeed, in Lemma 3.13 $K$ and $r_{0}$ depend on the minimizer, and consequently $c$ in Theorem 3.11 also inherits this dependence. At the same time, Lemma 4.2 requires $g$ to be $q$-growing, which is not required in Lemma 3.13.
@@ 4.1. Absence of holes in minimizers
Let $g$ be $\mathscr{R}$-admissible, 00-growing, infinitesimal and symmetric and let $G$ be $\mathscr{G}$-admissible and coercive. There exists $\bar{m}>0$, depending on $n$, $\lambda$, $g$ and $G$, such that, for all $m\in(0,\bar{m})$, every minimizer $E$ of $\mathscr{F}^{\lambda}$ with $|E|=m$ has the following property. There is no set $F\subset\mathbb{R}^{n}\setminus(H\cup E)$ with $|F|>0$ such that
@@
Let $g$ be $\mathscr{R}$-admissible, $q$-growing, infinitesimal and symmetric and let $G$ be $\mathscr{G}$-admissible and coercive. There exist $\bar{m}>0$ and $\bar{T}>0$, depending on $n$, $\lambda$, $g$, $G$, $q$ such that, for all $m\in(0,\bar{m})$, every volume constrained minimizer $E$ of $\mathscr{F}^{\lambda}$ with $|E|=m$ satisfies
@@
We remark that Lemma 4.8 is a stronger result than Lemma 3.12. Indeed, the bound in Lemma 4.8 does not depend on the minimizer. At the same time, Lemma 3.12 does not require that $g$ is $q$-growing and $G$ is coercive.
@@ 5. Generalized minimizers
Let $g$ be $\mathscr{R}$-admissible and $q$-growing and let $G$ be $\mathscr{G}$-admissible. For every $\varepsilon_{1},\varepsilon_{2}>0$ there exists a minimizer of ~$\tilde{\mathscr{F}}^{\lambda}_{\varepsilon}$ in the class
@@
Let $E\subset\mathbb{R}^{n}\setminus H$ be a set of finite perimeter. Let $g$ be $\mathscr{R}$-admissible and $q$-growing and let $G$ be $\mathscr{G}$-admissible. If $\alpha>1$, then
@@
$\sigma(b^{-})\leq\frac{1}{4}n|B^{\lambda}|^{\frac{1}{n}}m^{\frac{n-1}{n}}.$
@@
Let $g$ be $\mathscr{R}$-admissible and $q$-growing and let $G$ be $\mathscr{G}$-admissible. Let $\varepsilon_{1},\varepsilon_{2}>0$. For every $m\in(0,\infty)$ there exist $R>0$ and $\bar{h}\in\mathbb{N}$, depending on $n$, $\lambda$, $m$, $\varepsilon_{1}$, $\varepsilon_{2}$, $g$, $G$ and $q$, such that
@@
$m_{1}\geq\left(\frac{n|B^{\lambda}|^{\frac{n+1}{n}}}{4K}\right)^{n}.$
@@
$\alpha=\left(\frac{|B^{\lambda}|}{|B^{\lambda}|-\sum_{j>h^{\prime}}m_{j}}\right)^{\frac{1}{n}}\leq 1+c_{1}\sum_{j=h^{\prime}+1}^{h}m_{j},$
@@
Note that ~$\tilde{\mathscr{F}}^{\lambda}(\tilde{F})=\bar{\varepsilon}^{n-1}\tilde{\mathscr{F}}_{\bar{\varepsilon}}^{\lambda}(F)$ and that, if $g$ is $q$-growing, also $\tilde{g}$ is $q$-growing (see Remark 5.2). By Theorem 5.3 there exists $E\subset\mathbb{R}^{n}\setminus H$ with $|E|=|B^{\lambda}|$ which minimizes ~$\tilde{\mathscr{F}}^{\lambda}_{\bar{\varepsilon}}$. Then the set $\tilde{E}:=\bar{\varepsilon}E$ minimizes ~$\tilde{\mathscr{F}}^{\lambda}$ among sets with volume $m$ and Theorem 1.3 easily follows. ∎
Acknowledgements
The author is member of INdAM - GNAMPA. The author is grateful to Marco Pozzetta for many suggestions and for stimulating discussions. The author also thanks the referee for all valuable comments helping to concretely improve exposition of the results.